OpenStudy (anonymous):

A carpet manufacturer has received a special order from a contractor for precision-cut carpet. The carpet is rectangular, and the dimensions must be exactly 4+sqrt 7 yards wide by 5+sqrt 7 yards long. To know what the price of the carpet is, the salesman must know the total area of the carpet to be prepared for the contractor. What is the area of the carpet? (Recall that the area of a rectangle is the width multiplied by the length). A.27 square yards B. 27 + √7square yards C. 27 + 9√7 square yards D. 29 + 9√7 square yards.

OpenStudy (anonymous):

Area of rec: Length x Breath = (4 + sqrt 7)(5 + sqrt 7) = (4)(5) + (4)(sqrt 7) + (5)(sqrt 7) + (sqrt 7)(sqrt 7) = ?

OpenStudy (johnweldon1993):

and remember sqrt 7 times sqrt 7 = 7 so 4 times 5 = 20 then add the 7 from the sqrt 7 times sqrt 7..so 27 and add the 4sqrt7 + 5sqrt7 (when there is a sqrt7 on both of them...you just add the outside numbers and leave the sqrt7 alone...so 4sqrt7 + 5sqrt7 = 9sqrt7 so altogether you have 27 + 9sqrt7

OpenStudy (anonymous):

Thanks!!